### v1

# Bidirectional Nonlinear Optical Tomography: Unbiased Characterization of Off- and On-Chip Coupling Efficiencies

Abstract
Accurate evaluation of nonlinear photonic integrated circuits requires separating input and output coupling efficiencies (i.e., $\eta_{1}$ and $\eta_{2}$), yet the conventional linear-transmission calibration method recovers only their product (i.e., $\eta_{1}\eta_{2}$) and therefore introduces systematic bias when inferring on-chip performance from off-chip data. We present bidirectional nonlinear optical tomography (BNOT), a direction-aware metrology that uses forward and backward pumping of complementary nonlinear probes, with process-appropriate detection, to break the “degeneracy” of $\eta_{1}\eta_{2}$ and estimate individual interface efficiencies with tight confidence intervals. The method links off-chip measurements to on-chip quantities through a compact observation model that explicitly incorporates pump fluctuations and detector noise, and it frames efficiency extraction as a joint constrained optimization. Monte Carlo studies show unbiased convergence of the estimated efficiencies to ground truth with low error across realistic operating regimes. Using these efficiency estimates to reconstruct on-chip nonlinear figures of merit yields distributions centered on the true values with reduced variance, whereas conventional “degenerate” calibration is biased and can substantially misestimate on-chip performance. BNOT is hardware-compatible and platform-agnostic, and it provides unbiased unbiased characterization of off- and on-chip coupling efficiencies across nonlinear processes, enabling reproducible, coupling-resolved benchmarking for scalable systems in quantum optics, frequency conversion, and precision metrology.
[…]
Nonlinear photonic integrated circuits (PICs) promise a compact wafer-scale platform that transforms passive chips into active nonlinear engines for frequency conversion and generation of quantum states [1, 2]. Across emerging platforms such as lithium niobate (LiN), silicon nitride (SiN), aluminum nitride, and gallium arsenide, canonical processes, including spontaneous parametric down conversions (SPDC) [3, 4, 5], four-wave mixing (FWM) [6, 7, 8, 9, 10, 11, 12, 13], harmonic generation [14, 15, 16, 17], self-phase modulation (SPM) [18], optical parametric oscillation (OPO) [19, 20, 21, 22] and optical parametric amplification (OPA) [23, 3, 16], are now accessible on chip, enabling applications in spectroscopy, sensing, and quantum information science [3, 6, 7, 8, 9, 14, 15, 19, 20, 24]. These processes are carried out in diverse PIC structures (Fig. 1(a)), ranging from waveguides to a variety of microring resonators, which provide tailored dispersion and mode confinement. Multimode and microcomb architectures further scale brightness and mode count with favorable hardware efficiency [25], positioning nonlinear PICs as strong candidates for continuous-variable (CV) quantum computing [7, 26, 27] and classical optical computing [28, 29, 30, 31].
A major challenge in nonlinear PICs is the lack of bias-free performance estimation, especially as nonlinear interaction strength increases. Because experiments typically rely on off-chip lasers and detectors, the reported metrics involve two coupling interfaces whose efficiencies $(\eta_{1},\eta_{2})$ are often unequal due to mode mismatch, polarization sensitivity, fabrication imperfections, or differing coupling schemes (e.g., edge coupler on one side and grating coupler on the other). Conventional linear-transmission measurements yield only the product $\propto\eta_{1}\eta_{2}$, obscuring which interface dominates the loss and leading to the “degenerate” estimate $(\eta_{1}\eta_{2})^{1/2}$, a source of systematic bias in evaluating on-chip performance [32, 33, 34, 32]. An example is chip-based second harmonic generation (SHG) measured off chip, where the detected efficiency depends asymmetrically on the two interfaces. This can be expressed as $f_{AS}(\eta_{1},\eta_{2})$, satisfying $f_{AS}(\eta_{1},\eta_{2})\neq f_{AS}(\eta_{2},\eta_{1})$ and scaling as $P_{\text{SHG}}/P_{\text{in}}^{2}\propto\eta_{\text{in}}^{2}\eta_{\text{out}}$ [35]. Conventional calibration assumes $\eta_{\text{in}}=\eta_{\text{out}}=(\eta_{1}\eta_{2})^{1/2}$, giving $P_{\text{SHG}}/P_{\text{in}}^{2}\propto(\eta_{1}\eta_{2})^{3/2}$ (curve (i) in Fig. 1(b)), whereas the actual response depends on direction: forward pumping ($\eta_{\text{in}}=\eta_{1}$, $\eta_{\text{out}}=\eta_{2}$) yields $\propto\eta_{1}^{2}\eta_{2}$ (curve (ii)), and backward pumping ($\eta_{\text{in}}=\eta_{2}$, $\eta_{\text{out}}=\eta_{1}$) yields $\propto\eta_{2}^{2}\eta_{1}$ (curve (iii)). This intrinsic asymmetry ultimately explains the systematic deviation observed under symmetric calibration. Wang et al. [36] used forward and backward SHG as symmetric and asymmetric references to attribute performance differences to the different input/output coupling efficiencies. However, such approaches still assume equal efficiencies in the symmetric case and cannot uniquely resolve individual efficiencies.
Among nonlinear processes, squeezed light provides another stringent benchmark for on-chip device performance. Bulk-optics squeezing experiment have reached $15\penalty 10000\ \text{dB}$ [37], while envisioned applications require near-$10\penalty 10000\ \text{dB}$ for CV fault-tolerant quantum computing [26], gravitational-wave detectors [38], and Gottesman-Kitaev-Preskill (GKP) sources [39]. Integrated platforms have advanced from $1.7\penalty 10000\ \text{dB}$ demonstrations [10] to multi-decibel squeezing in LiN [7], cavity optomechanics [40], and foundry-compatible SiN [41, 12], with further gains in nanophotonic molecules and microcombs [42, 11]. The measured squeezing levels reported (off-chip) now exceed $4.9\penalty 10000\ \text{dB}$ in PPLN waveguides, $3.5\penalty 10000\ \text{dB}$ in Kerr microrings, $5.6\penalty 10000\ \text{dB}$ in microcombs [3, 6, 9], and attain $3.1\penalty 10000\ \text{dB}$ in wafer-scale integration [12]. As measured values increase and target on-chip squeezing thresholds approach $10-15\penalty 10000\ \text{dB}$, an accurate estimation of interface efficiencies $(\eta_{1},\eta_{2})$ is becoming even more critical [43, 44, 6, 45].
We introduce bidirectional nonlinear optical tomography (BNOT), a directionally-sensitive metrology that combines forward- and backward-pumped nonlinear probes to break the $\eta_{1}\eta_{2}$ “degeneracy” inherent to linear transmission. BNOT enables direct estimation of the individual interface coupling efficiencies with confidence intervals that are significantly narrower than those obtained from conventional linear calibration.
Figure 2: Histograms of Monte Carlo (MC) simulations with the ground-truth interface efficiencies in Eq. (5) and parameters in Tab. 1. (a) Simulated off-chip squeezing and (b) SHG efficiency. (c) BNOT estimates of $\hat{\eta}_{1}$ (blue) and $\hat{\eta}_{2}$ (red), with shaded regions showing 95 % CIs (2.5–97.5 percentiles).
I Concept
Our BNOT methodology leverages the reversibility of two nonlinear processes within the same photonic device by pumping it in opposite directions. To benchmark its performance, we consider on-chip squeezed-light generation and SHG in an exemplary platform: periodically poled LiN (PPLN) waveguides. The calibration scheme is illustrated in Fig. 1(c). A pump field at $2\omega$ is injected into the PPLN waveguide through the left interface in the forward direction (left-to-right), while a pump at $\omega$ is simultaneously launched through the right interface in the backward direction (right-to-left). The corresponding interface coupling efficiencies are denoted $\eta_{1}^{(2\omega)}$, $\eta_{2}^{(2\omega)}$, $\eta_{1}^{(\omega)}$, and $\eta_{2}^{(\omega)}$, where the superscript denotes the pump frequency. This bidirectional pumping exploits the reversibility of the nonlinear interactions, enabling both degenerate SPDC and SHG to occur concurrently. The corresponding output fields are measured at their respective ports using balanced homodyne detection and direct detection.
To relate on- and off-chip performances in squeezing and SHG, we define two estimation variables, $x_{1}$ and $x_{2}$, corresponding to the left- and right-coupling efficiencies of the PPLN waveguide. Notably, these are not the ground-truth efficiencies: $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ but the tunable parameters in our model used to recover them. The off-chip squeezing level (in dB) and SHG efficiency (in W-1m-2) are expressed as
[…]
$\mathcal{E}_{\text{OFF}}\,(x_{1},x_{2})$ $=x_{1}\,\mathcal{E}_{\text{ON}}(x_{2})+\xi_{\text{shg}},$
where
$\mathcal{S}_{\text{ON}}\,(x_{1})$ $=2\sqrt{x_{1}\,\alpha\,P_{2\omega}\,\varepsilon_{1}}\,\mathrm{sinc}\left(\Delta k\,L/2\right)\;\;\;\;\text{and}$ (2)
$\mathcal{E}_{\text{ON}}\,(x_{2})$ $=x_{2}^{2}\,\varepsilon_{2}^{2}\,\alpha\,\mathrm{sinc}^{2}\left(\Delta k\,L/2\right)$
denote the analytical formula for the on-chip squeezing [46] and the SHG efficiency [17]. Here $P_{2\omega}$ denotes the non-depleted off-chip pump power at $2\omega$ for forward pumping, $\varepsilon_{1}\sim U(1-\delta_{1},1+\delta_{1})$ and $\varepsilon_{2}\sim U(1-\delta_{2},1+\delta_{2})$ are the uniform fluctuations of the corresponding pump powers, $\delta_{1}$ and $\delta_{2}$ are the half of the corresponding fluctuation ranges, $\xi_{\text{sqz}}\sim\mathcal{N}(0,\sigma^{2}_{\text{sqz}})$ and $\xi_{\text{shg}}\sim\mathcal{N}(0,\sigma^{2}_{\text{shg}})$ represent additive Gaussian noise in the squeezing and SHG measurements (e.g., instability of the measurement system), and $\alpha=8\pi^{2}d_{\mathrm{eff}}^{2}L^{2}/\varepsilon_{0}cn_{2\omega}n_{\omega}^{2}\lambda_{\omega}^{2}A$. The parameter $d_{\mathrm{eff}}$ is the effective nonlinear coefficient under quasi-phase-matching, $L$ is the poling length, $\varepsilon_{0}$ is the vacuum permittivity, $c$ is the speed of light, $n_{2\omega}$ and $n_{\omega}$ are the refractive indices of lithium niobate at $2\omega$ and $\omega$, respectively, and $\Delta k$ and $A$ denote the wave-vector mismatch and waveguide mode area (i.e., $\lambda_{\omega}=2\pi c/\omega$).
Figure 3: (a) Total root-mean-square error (RMSE) map of efficiency estimation using BNOT, $\sqrt{e_{1}^{2}+e_{2}^{2}}$, where $e_{1}$ and $e_{2}$ are the RMSEs of $\eta^{(2\omega)}_{1}$ and $\eta^{(\omega)}_{2}$, respectively. The red dashed contour marks the $0.01$ boundary. Colored stars indicate the parameter sets used for MC simulations. (b) RMSE of $\eta^{(2\omega)}_{1}$ and (c) RMSE of $\eta^{(\omega)}_{2}$ as functions of the number of independent trials, with colors matching the parameter sets in (a).
The conventional calibration approach based on linear transmission cannot distinguish between left-to-right and right-to-left pumping [32, 33, 34, 32]. As a result, it enforces a “degenerate”-interface assumption, estimating the efficiencies as
$\hat{\eta}_{1,\text{C}}=\hat{\eta}_{2,\text{C}}$ $=\left(\eta_{2}^{(\omega)}\eta_{1}^{(\omega)}\right)^{1/2}\;\;\;\;\;\,\text{at $\omega$},$ (3)
$=\left(\eta_{2}^{(2\omega)}\eta_{1}^{(2\omega)}\right)^{1/2}\;\;\;\text{at $2\omega$}.$
However, this assumption is unreliable since $\eta^{(2\omega)}_{1}$, $\eta^{(2\omega)}_{2}$, $\eta^{(\omega)}_{1}$, and $\eta^{(\omega)}_{2}$ can all differ substantially. This “degeneracy” constraint introduces significant uncertainty and yields a biased estimator of on-chip performance. In particular, when $\mathcal{S}_{\text{ON}}$ or $\mathcal{E}_{\text{ON}}$ is inferred by back-calculating from measured off-chip quantities $\mathcal{S}_{\text{OFF}}$ and $\mathcal{E}_{\text{OFF}}$ (Eq. (1)), the use of $x_{2}=\hat{\eta}_{2,\text{C}}$ under conventional calibration introduces systematic error.
Figure 4: MC simulations of squeezing and SHG efficiency. Histograms of (a) off-chip squeezing and (b) SHG efficiency. The estimated (c) on-chip squeezing and (d) SHG efficiency. Black dashed lines mark the ground-truth of on-chip squeezing and SHG efficiency: $\mathcal{S}_{\text{ON}}(\eta^{(2\omega)}_{1})$ and $\mathcal{E}_{\text{ON}}(\eta^{(\omega)}_{2})$. In (c) and (d), the brown histograms represent conventional calibration assuming “degenerate” coupling efficiencies $\hat{\eta}_{1,C}=\hat{\eta}_{2,C}=0.784$ (pumping at $\omega$), while the green histograms show the estimates of BNOT.
To overcome this challenge, we introduce an optimization-based estimation method to systematically search the parameter space for optimal estimates $\hat{\eta}_{1}$ and $\hat{\eta}_{2}$. Specifically, the estimation of the two efficiencies is defined as the solution of the following optimization:
$\hat{\eta}_{1,\text{B}}\;,\;\hat{\eta}_{2,\text{B}}=\underset{{x_{1},\,x_{2}\,\in\left(0,1\right)}}{\text{argmin}}\left\{e_{\text{sqz}}^{2}(x_{1},x_{2})+e_{\text{shg}}^{2}(x_{1},x_{2})\right\},$ (4)
where $e^{2}_{\text{sqz}}(x_{1},x_{2})\equiv\left|1-\mathcal{S}_{\text{OFF}}(x_{1},x_{2})/\mathcal{S}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$ and $e^{2}_{\text{shg}}(x_{1},x_{2})\equiv\left|1-\mathcal{E}_{\text{OFF}}(x_{1},x_{2})/\mathcal{E}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$. In words, the coupling efficiencies $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ are estimated by determining the pair of values $\hat{\eta}_{1,\text{B}},\hat{\eta}_{2,\text{B}}$ that minimizes the equally weighted mean square error rate for the squeezing and SHG efficiency estimation. To evaluate the estimation outcome with BNOT, we consider the ground-truth interface efficiencies:
[…]
and adopt the Monte Carlo (MC) simulation, using the physical parameters of the PPLN waveguide in Tab. 1. Fig. 2(c) shows histograms of $\hat{\eta}_{1,\text{B}}$ and $\hat{\eta}_{2,\text{B}}$ obtained from MC simulations. The close agreement with the ground-truth values confirms the unbiasedness of our estimators. However, with the conventional calibration approach, the estimation is $\hat{\eta}_{1,\text{C}}=\hat{\eta}_{2,\text{C}}=0.784$ for pump at $\omega$ or $0.741$ for pump at $2\omega$, both of which deviate from the pair of ground-truth efficiencies: $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ (black dashed vertical lines in Fig. 2(c)) that are relevant to squeezing and SHG experiments.
To assess general ground-truth efficiencies, we scanned the full range of $\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2}\in\left[0,1\right]$. For each pair $(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})$, our estimator is applied using MC simulations, with the total root-mean-square error (RMSE) of $\hat{\eta}_{1,\text{B}}$ and $\hat{\eta}_{2,\text{B}}$, $\sqrt{e_{1}^{2}+e_{2}^{2}}$, shown in Fig. 3(a). Here, $e_{1}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{1,\text{B}}-\eta_{1}^{(2\omega)})^{2}\right]}$ and $e_{2}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{2,\text{B}}-\eta_{2}^{(\omega)})^{2}\right]}$ denote the RMSEs of $\hat{\eta}_{1,\text{B}}$ and $\hat{\eta}_{2,\text{B}}$, and $\mathbb{E}[\cdot]$ denotes the arithmetic mean in simulation trials. Five representative pairs of $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$, indicated by the colored stars in Fig. 3(a). The corresponding RMSEs, $e_{1}$ and $e_{2}$, are shown in Fig. 3(b,c).
From Fig. 3(a), achieving a low RMSE imposes a much stricter requirement on the estimation of $\eta_{2}^{(\omega)}$ than on $\eta_{1}^{(2\omega)}$. For instance, the total estimation error remains $\leq 0.01$ (red dashed curve) only when $\eta_{1}^{(2\omega)}\geq 0.3$ and $\eta_{2}^{(\omega)}\geq 0.85$. This asymmetry arises because the detected squeezing depends nonlinearly on the output efficiency $\eta_{2}^{(\omega)}$, which enters exponentially in the measured quadrature variance. As a result, small deviations in $\eta_{2}^{(\omega)}$ strongly affect the observed squeezing. In contrast, SHG depends linearly on $\eta_{2}^{(\omega)}$ and quadratically on $\eta_{1}^{(2\omega)}$, distributing the effect of measurement noise $\xi_{\text{shg}}$ more symmetrically across both interfaces.
Using the ground-truth efficiencies defined in Eq. (5), we estimate $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ with both the conventional calibration and BNOT methods. These values are then used to back-calculate the on-chip squeezing and SHG efficiency via the noiseless relations in Eq. (1). Based on the simulated off-chip data in Fig. 4(a,b), the reconstructed on-chip performances are shown in Fig. 4(c,d) with the associated ground truth (black dashed vertical lines). As expected, the BNOT estimator remains unbiased, whereas the conventional calibration exhibits systematic bias. Because BNOT jointly fits the nonlinear processes, the estimation of $\eta_{2}^{(\omega)}$, which strongly affects off-chip squeezing, directly constrains $\mathcal{S}_{\text{ON}}$ through a shared likelihood function. This self-consistency reduces both bias and variance in the reconstructed on-chip squeezing and SHG efficiency. Conversely, the conventional approach treats the two calibration steps independently, leading to systematically decorrelated estimates. This is most pivotal in high-squeezing regimes (e.g., $\mathcal{S}_{\text{OFF}}\sim 15\penalty 10000\ \text{dB}$), where BNOT’s unbiased posteriors for $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$ translate directly into precise interface-efficiency requirements and the corresponding $\mathcal{S}_{\text{ON}}$ needed to reach the target (see Appendix).
II Outlook
In summary, BNOT provides a platform-agnostic approach to calibrate asymmetric coupling efficiencies in nonlinear PICs. BNOT works with SHG and squeezed-light measurements, so it can be deployed on existing setups without extra hardware to benchmark chip-to-fiber and inter-stage interfaces on the chip. An exciting near-term application of BNOT can be in high-device-yield nonlinear PICs experiments, such as recent demonstrations of wafer-scale on-chip multi-harmonic generation [15] and on-chip squeezing [12]. Beyond our exemplary investigation of on-chip squeezing and SHG, BNOT provides an unbiased calibration approach for other on-chip nonlinear phenomena [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 19, 20, 21, 22, 23, 3, 16].
Looking ahead, an intriguing direction may be harnessing BNOT in adaptive control and self-calibrating architectures. By combining real-time forward/backward measurements with Bayesian [47, 48, 49, 50] or machine-learning estimators [51], one can automate coupling-loss compensation across large photonic arrays and dynamically track degradation or drift. Such closed-loop implementations may convert coupling characterization from a post-hoc diagnostic into an active subsystem of photonic hardware, supporting reproducible, system-level benchmarks as nonlinear PICs transition from laboratory demonstrations to scalable quantum and classical technologies.
III Acknowledgments
The authors thank Shi-Yuan Ma, Mengjie Yu, Avik Dutt, and Sri Krishna Vadlamani for fruitful discussions. This work was supported by funding from the DARPA INSPIRED program.
IV Appendix
MC simulation parameters $d_{\text{eff}}$ $19.5$ pm/V Thin-film LiN effective nonlinear coefficient $L$ 66 mm  Length of PPLN waveguide AA $1.4$ µm2 Cross-sectional area of PPLN waveguide $\omega$ $193$ THz Pumping frequency (or $2\omega$)  (or $386$ THz)   at $1550$ nm (or $775$ nm) $n_{\omega}$ $2.18$  Refractive index of LiN (or $n_{2\omega}$) $2.14$  at $\omega$ (or $2\omega$) $\sigma_{\text{sqz}}$ $0.1\penalty 10000\ \text{dB}$  Squeezing (or SHG) (or $\sigma_{\text{shg}}$)    (or $20\penalty 10000\ \text{\%/W/cm${}^{2}$}$)   measurement noise STD $P_{2\omega}$ $100$ mW   Off-chip mean pump (or $P_{\omega}$) (or $10$ mW) power in squeezing (or SHG) experiment Off-chip pump power $\delta_{1}$ $0.03$ half of the fluctuation (or $\delta_{2}$) (or $0.05$) rate in squeezing (or SHG) experiment $\Delta k$ 0 m-1 Wave-vector mismatch
Table 1: MC simulation parameters of PPLN waveguide in squeezing and SHG experiments. STD: standard deviation.
IV.1 Efficiency thresholds for off-chip squeezing
Figure 5: (a) On-chip squeezing versus $\eta_{1}^{(2\omega)}$. (b) Off-chip squeezing versus $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$. The orange horizontal line in (a) and curve in (b) denote the on-chip and off-chip $15$-dB squeezing threshold.
Fig. 5 summarizes the dependence of on- and off-chip squeezing on interface efficiencies. Fig. 5(a) shows the on-chip squeezing $\mathcal{S}_{\text{ON}}$ as a function of the coupling efficiency $\eta^{(2\omega)}_{1}$, while Fig. 5(b) maps the off-chip squeezing $\mathcal{S}_{\text{OFF}}$ as a joint function of $\eta^{(2\omega)}_{1}$ and $\eta^{(\omega)}_{2}$. The orange lines in both panels mark the squeezing contour $15\penalty 10000\ \text{dB}$, a benchmark frequently cited in advanced quantum technologies.
On-chip squeezing levels approaching $15\penalty 10000\ \text{dB}$ are particularly relevant for applications such as CV quantum error correction and precision quantum metrology [26, 39]. As shown in Fig. 5(a), achieving this on-chip benchmark requires an input coupling efficiency $\eta^{(2\omega)}_{1}\gtrsim 0.65$, which is attainable using current photonic integration technologies. For example, edge-coupled LiN modulators have demonstrated per-interface coupling efficiencies below $0.5\penalty 10000\ \text{dB}$ (corresponding to efficiencies above $90\penalty 10000\ \%$) at $1550\penalty 10000\ \text{nm}$ [52], indicating that such coupling efficiencies are experimentally feasible.
Although high on-chip squeezing is feasible, maintaining this performance off chip imposes stricter requirements. For example, to retain the off-chip squeezing level $\mathcal{S}_{\text{OFF}}\gtrsim 15\penalty 10000\ \text{dB}$, the output interface efficiency must satisfy $\eta^{(\omega)}_{2}\gtrsim 0.97$, illustrated in Fig. 5(b).