### v2

# Bidirectional Nonlinear Optical Tomography: Unbiased Characterization of Off- and On-Chip Coupling Efficiencies

Abstract
Accurate evaluation of nonlinear photonic integrated circuits requires separating input and output coupling efficiencies (i.e., $\eta_{1}$ and $\eta_{2}$), yet the conventional linear-transmission calibration method recovers only their product (i.e., $\eta_{1}\,\eta_{2}$) and therefore introduces systematic bias when inferring on-chip performance from off-chip data. We present bidirectional nonlinear optical tomography (BNOT), a direction-aware metrology that uses forward and backward pumping of complementary nonlinear probes, with process-appropriate detection, to break the “degeneracy” of $\eta_{1}\,\eta_{2}$ and estimate individual interface efficiencies with tight confidence intervals. The method links off-chip measurements to on-chip quantities through a compact observation model that explicitly incorporates pump fluctuations and detector noise, and it frames efficiency extraction as a joint constrained optimization. Monte Carlo studies show unbiased convergence of the estimated efficiencies to ground truth with low error across realistic operating regimes. Using these efficiency estimates to reconstruct on-chip nonlinear figures of merit yields distributions centered on the true values with reduced variance, whereas conventional “degenerate” calibration is biased and can substantially misestimate on-chip performance. BNOT is hardware-compatible and platform-agnostic, and provides unbiased characterization of off- and on-chip coupling efficiencies across nonlinear processes, enabling reproducible, coupling-resolved benchmarking for scalable systems in quantum optics, frequency conversion, and precision metrology.
[…]
I Introduction
Nonlinear photonic integrated circuits (PICs) promise a compact wafer-scale platform that transforms passive chips into active nonlinear engines for frequency conversion and generation of quantum states [1, 2]. Across emerging platforms such as lithium niobate (LiN), silicon nitride (SiN), aluminum nitride, and gallium arsenide, canonical processes, including spontaneous parametric down conversion (SPDC) [3, 4, 5], four-wave mixing (FWM) [6, 7, 8, 9, 10, 11, 12, 13], harmonic generation [14, 15, 16, 17], self-phase modulation (SPM) [18], optical parametric oscillation (OPO) [19, 20, 21, 22] and optical parametric amplification (OPA) [23, 3, 16], are now accessible on chip, enabling applications in spectroscopy, sensing, and quantum information science [3, 6, 7, 8, 9, 14, 15, 19, 20, 24]. These processes are carried out in diverse PIC structures (Fig. 1(a)), ranging from waveguides to a variety of microring resonators, which provide tailored dispersion and mode confinement. Multimode and microcomb architectures further scale brightness and mode count with favorable hardware efficiency [25], positioning nonlinear PICs as strong candidates for continuous-variable (CV) quantum computing [7, 26, 27] and classical optical computing [28, 29, 30, 31].
A major challenge in nonlinear PICs is the lack of bias-free performance estimation, especially as nonlinear interaction strength increases. Because experiments typically rely on off-chip lasers and detectors, the reported metrics involve two coupling interfaces whose efficiencies $(\eta_{1},\eta_{2})$ are often unequal due to mode mismatch, polarization sensitivity, fabrication imperfections, or differing coupling schemes (e.g., edge coupler on one side and grating coupler on the other). Conventional linear-transmission measurements yield only the product $\propto\eta_{1}\,\eta_{2}$, obscuring which interface dominates the loss and leading to the “degenerate” estimate $(\eta_{1}\,\eta_{2})^{1/2}$, a source of systematic bias in evaluating on-chip performance [32, 33, 34, 32]. An example is chip-based second harmonic generation (SHG) measured off chip, where the detected efficiency depends asymmetrically on the two interfaces. This can be expressed as $f_{AS}(\eta_{1},\eta_{2})$, satisfying $f_{AS}(\eta_{1},\eta_{2})\neq f_{AS}(\eta_{2},\eta_{1})$ and scaling as $P_{\text{SHG}}/P_{\text{in}}^{2}\propto\eta_{\text{in}}^{2}\eta_{\text{out}}$ [35]. Conventional calibration assumes $\eta_{\text{in}}=\eta_{\text{out}}=(\eta_{1}\,\eta_{2})^{1/2}$, giving $P_{\text{SHG}}/P_{\text{in}}^{2}\propto(\eta_{1}\,\eta_{2})^{3/2}$ (curve (i) in Fig. 1(b)), while the actual response depends on direction: forward pumping ($\eta_{\text{in}}=\eta_{1}$, $\eta_{\text{out}}=\eta_{2}$) yields $\propto\eta_{1}^{2}\,\eta_{2}$ (curve (ii)), and backward pumping ($\eta_{\text{in}}=\eta_{2}$, $\eta_{\text{out}}=\eta_{1}$) yields $\propto\eta_{2}^{2}\,\eta_{1}$ (curve (iii)). This intrinsic asymmetry ultimately explains the systematic deviation observed under symmetric calibration. Wang et al. [36] used forward and backward SHG as symmetric and asymmetric references to attribute performance differences to the different input/output coupling efficiencies. However, such approaches still assume equal efficiencies in the symmetric case and cannot uniquely resolve individual efficiencies.
Among nonlinear processes, squeezed light provides another stringent benchmark for on-chip device performance. Bulk-optics squeezing experiments have reached $15\penalty 10000\ \text{dB}$ [37], while envisioned applications require near-$10\penalty 10000\ \text{dB}$ for CV fault-tolerant quantum computing [26], gravitational-wave detectors [38], and Gottesman-Kitaev-Preskill (GKP) sources [39]. Integrated platforms have advanced from $1.7\penalty 10000\ \text{dB}$ demonstrations [10] to multi-decibel squeezing in LiN [7], cavity optomechanics [40], and foundry-compatible SiN [41, 12], with further gains in nanophotonic molecules and microcombs [42, 11]. The measured squeezing levels reported (off-chip) now exceed $4.9\penalty 10000\ \text{dB}$ in PPLN waveguides, $3.5\penalty 10000\ \text{dB}$ in Kerr microrings, $5.6\penalty 10000\ \text{dB}$ in microcombs [3, 6, 9], and attain $3.1\penalty 10000\ \text{dB}$ in wafer-scale integration [12]. As measured values increase and target chip squeezing thresholds approach $10-15\penalty 10000\ \text{dB}$, an accurate estimation of interface efficiencies $(\eta_{1},\eta_{2})$ becomes even more critical [43, 44, 6, 45].
We introduce bidirectional nonlinear optical tomography (BNOT), a directionally-sensitive metrology that combines forward- and backward-pumped nonlinear probes to break the $\eta_{1}\,\eta_{2}$ “degeneracy” inherent to linear transmission. BNOT enables direct estimation of the individual interface coupling efficiencies with confidence intervals that are significantly narrower than those obtained from conventional linear calibration.
Figure 2: Histograms of Monte Carlo (MC) simulations with the ground-truth interface efficiencies in Eq. (4) and parameters in Tab. 1. (a) Simulated off-chip squeezing and (b) SHG efficiency. (c) BNOT estimates of $\hat{\eta}_{1,\,\text{B}}$ (blue) and $\hat{\eta}_{2,\,\text{B}}$ (red), with shaded regions showing 95 % confidence intervals (2.5–97.5 percentiles).
II Concept
Our BNOT methodology leverages the reversibility of two nonlinear processes within the same photonic device by pumping it in opposite directions. To benchmark its performance, we consider on-chip squeezed-light generation and SHG in an exemplary platform: periodically poled LiN (PPLN) waveguides. The calibration scheme is illustrated in Fig. 1(c). A pump field at $2\omega$ is injected into the PPLN waveguide through the left interface in the forward direction (left-to-right), while a pump at $\omega$ is simultaneously launched through the right interface in the backward direction (right-to-left). The corresponding interface coupling efficiencies are denoted $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$, where the superscript denotes the pump frequency. This bidirectional pumping exploits the reversibility of the nonlinear interactions, enabling both degenerate SPDC and SHG to occur concurrently. The corresponding output fields are measured at their respective ports using balanced homodyne detection and direct detection.
To relate on- and off-chip performances in squeezing and SHG, we define two estimation variables, $x_{1}$ and $x_{2}$, corresponding to the left- and right-coupling efficiencies of the PPLN waveguide. In particular, these are not the ground-truth efficiencies: $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ but the tunable parameters in our model used to recover them. The off-chip squeezing level (in dB) and SHG efficiency (in W-1m-2) are expressed as
[…]
$\mathcal{E}_{\text{OFF}}\,(x_{1},x_{2})$ $=x_{1}\,\mathcal{E}^{(\varepsilon_{2}\,,\,\Delta k)}_{\text{ON}}(x_{2})+\xi_{\text{shg}},$
where $\mathcal{S}^{(\varepsilon_{1},\,\Delta k)}_{\text{ON}}$ and $\mathcal{E}^{(\varepsilon_{2},\,\Delta k)}_{\text{ON}}$ denote the analytical models for the on-chip squeezing [46] and SHG efficiency [17], respectively, with their derivations provided in App. A. The parameters $\varepsilon_{1}\sim U(1-\delta_{1},1+\delta_{1})$ and $\varepsilon_{2}\sim U(1-\delta_{2},1+\delta_{2})$ describe the uniform fractional fluctuations in the pump powers, where $\delta_{1}$ and $\delta_{2}$ specify the corresponding half-widths of the fluctuation ranges. The wave-vector mismatch is modeled as $\Delta k\sim\mathcal{N}(0,\sigma_{\Delta k}^{2})$, with $\sigma_{\Delta k}$ representing the standard deviation arising from the temperature instability of the PPLN waveguide. Additional measurement noise in the squeezing and SHG signals is captured by $\xi_{\text{sqz}}\sim\mathcal{N}(0,\sigma_{\text{sqz}}^{2})$ and $\xi_{\text{shg}}\sim\mathcal{N}(0,\sigma_{\text{shg}}^{2})$, which account for residual technical fluctuations in the detection system.
Figure 3: (a) Total mean-square error (MSE) map of efficiency estimation using BNOT, $e_{1}^{2}+e_{2}^{2}$, where $e^{2}_{1}$ and $e^{2}_{2}$ are the estimator’s MSEs of $\eta^{(2\omega)}_{1}$ and $\eta^{(\omega)}_{2}$, respectively. The red dashed contour marks the 20-20 dB boundary. Colored stars indicate the parameter sets used for Monte Carlo simulations. (b) MSE of $\hat{\eta}_{1,\,\text{B}}$ and (c) MSE of $\hat{\eta}_{2,\,\text{B}}$ as functions of the number of independent trials, with colors matching the parameter sets in (a).
The conventional linear transmission-based calibration approach cannot distinguish between left-to-right and right-to-left pumping [32, 33, 34, 32]. As a result, it enforces a “degenerate”-interface assumption, estimating the efficiencies as
$\hat{\eta}_{1,\text{C}}=\hat{\eta}_{2,\text{C}}=\mathcal{N}\left(\sqrt{\eta_{1}^{(\omega)}\,\eta_{2}^{(\omega)}},\sigma_{\text{C}}^{2}\right),$ (2)
with the measurement standard deviation $\sigma_{\text{C}}$. Clearly, the mean of the estimate is biased by the ground-truth efficiencies that are relevant to squeezing and SHG experiments (i.e., $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$). This “degeneracy” constraint introduces significant uncertainty and yields a biased estimator of on-chip performance. In particular, when the on-chip $\mathcal{S}_{\text{ON}}$ or $\mathcal{E}_{\text{ON}}$ is inferred by back-calculating the measured quantities off the chip $\mathcal{S}_{\text{OFF}}$ and $\mathcal{E}_{\text{OFF}}$ (Eq. (1)), the conventional calibration procedure replaces $x_{2}$ with the estimated coupling efficiency $\hat{\eta}_{2,\text{C}}$, thus introducing a systematic bias in the inferred result.
Figure 4: MC simulations of squeezing and SHG efficiency. Histograms of (a) off-chip squeezing and (b) SHG efficiency. The estimated (c) on-chip squeezing and (d) SHG efficiency are shown. Black dashed lines mark the ground-truth of on-chip squeezing and SHG efficiency: $\mathcal{S}^{(\varepsilon_{1},\,\Delta k)}_{\text{ON}}$ and $\mathcal{E}^{(\varepsilon_{2},\,\Delta k)}_{\text{ON}}$. In (c) and (d), the brown histograms represent conventional calibration assuming “degenerate” coupling efficiencies, while the green histograms show the estimates of BNOT.
III Simulation Results
To overcome this challenge, we introduce an optimization-based estimation method to systematically search the parameter space for optimal estimates $\hat{\eta}_{1}$ and $\hat{\eta}_{2}$. Specifically, the estimation of the two efficiencies is defined as the solution of the following optimization:
$\hat{\eta}_{1,\,\text{B}}\;,\;\hat{\eta}_{2,\,\text{B}}=\underset{{x_{1},\,x_{2}\,\in\left(0,1\right)}}{\text{argmin}}\left\{e_{\text{sqz}}^{2}(x_{1},x_{2})+e_{\text{shg}}^{2}(x_{1},x_{2})\right\},$ (3)
where $e^{2}_{\text{sqz}}(x_{1},x_{2})\equiv\left|1-\mathcal{S}_{\text{OFF}}(x_{1},x_{2})/\mathcal{S}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$ and $e^{2}_{\text{shg}}(x_{1},x_{2})\equiv\left|1-\mathcal{E}_{\text{OFF}}(x_{1},x_{2})/\mathcal{E}_{\text{OFF}}(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})\right|^{2}$. In words, the coupling efficiencies $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ are estimated by determining the pair of values $\hat{\eta}_{1,\,\text{B}},\hat{\eta}_{2,\,\text{B}}$ that minimizes the equally weighted mean square error rate for the squeezing and SHG efficiency estimation.
We consider the ground-truth interface efficiencies:
[…]
and adopt the Monte Carlo (MC) simulation in Fig. 2(a, b), using the physical parameters of the PPLN waveguide in Tab. 1. Fig. 2(c) presents the histograms of the estimators $\hat{\eta}_{1,\,\text{B}}$ and $\hat{\eta}_{2,\,\text{B}}$ obtained from the MC simulations. Their close agreement with the ground-truth values confirms the unbiasedness of the proposed estimators. Moreover, Fig. 2(c) reveals that the estimation of $\eta_{2}^{(\omega)}$ exhibits a significantly narrower confidence interval than that of $\eta_{1}^{(2\omega)}$. This asymmetry arises because, in squeezing, the output efficiency $\eta_{2}^{(\omega)}$ directly sets the amount of vacuum noise mixed into the detected field, so small changes in $\eta_{2}^{(\omega)}$ cause large variations in the observed squeezing. In contrast, the SHG signal depends linearly on $\eta_{2}^{(\omega)}$ and quadratically on $\eta_{1}^{(2\omega)}$, which distributes the measurement noise $\xi_{\text{shg}}$ more evenly between the two interfaces.
To assess general ground-truth efficiencies, we scanned the full range of $\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2}\in\left[0,1\right]$. For each pair $(\eta^{(2\omega)}_{1},\eta^{(\omega)}_{2})$, our estimator is applied using MC simulations, with the overall mean-square error (MSE) of $\hat{\eta}_{1,\,\text{B}}$ and $\hat{\eta}_{2,\,\text{B}}$, $e_{1}^{2}+e_{2}^{2}$, in decibels, shown in Fig. 3(a) (the red dashed curve marks the MSE boundary of 20-20 dB). Here, $e_{1}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{1,\,\text{B}}-\eta_{1}^{(2\omega)})^{2}\right]}$ and $e_{2}\equiv\sqrt{\mathbb{E}\left[(\hat{\eta}_{2,\,\text{B}}-\eta_{2}^{(\omega)})^{2}\right]}$, where $\mathbb{E}[\cdot]$ denotes the arithmetic mean in simulation trials. Five representative pairs of $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$, indicated by the colored stars in Fig. 3(a). The corresponding MSEs, $e^{2}_{1}$ and $e^{2}_{2}$, are shown in Fig. 3(b,c).
Using the ground-truth efficiencies defined in Eq. (4), we estimate $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$ with both conventional calibration and BNOT methods. These values are then used to back-calculate the on-chip squeezing and SHG efficiency via the noiseless relations in Eq. (1). Based on the simulated off-chip data in Fig. 4(a,b), the reconstructed on-chip performances are shown in Fig. 4(c,d) with the associated ground truth (black dashed vertical lines). From the simulation result, the BNOT estimator remains unbiased, whereas the conventional calibration exhibits a systematic bias. Because BNOT jointly fits the nonlinear processes, the estimation of $\eta_{2}^{(\omega)}$, which strongly affects off-chip squeezing, directly constrains $\mathcal{S}_{\text{ON}}$ through a shared likelihood function. This self-consistency reduces both bias and variance in the reconstructed on-chip squeezing and SHG efficiency. In contrast, the conventional approach treats the two calibration steps independently, leading to systematically decorrelated estimates. This is most pivotal in high-squeezing regimes (e.g., $\mathcal{S}_{\text{OFF}}\sim 15\penalty 10000\ \text{dB}$), where BNOT’s unbiased posteriors for $(\eta_{1}^{(2\omega)},\eta_{2}^{(\omega)})$ translate directly into precise interface-efficiency requirements and the corresponding $\mathcal{S}_{\text{ON}}$ needed to reach the target (see App. B).
IV Conclusion and outlook
In summary, BNOT provides a platform-agnostic method to calibrate asymmetric coupling efficiencies in nonlinear PICs. As we prototypically demonstrated for SHG and squeezed-light measurements, BNOT integrates seamlessly with existing experimental setups and enables accurate benchmarking of both chip-to-fiber interfaces and inter-stage connections. This capability is particularly valuable for high–device-yield nonlinear PIC platforms, including recent wafer-scale demonstrations of on-chip multi-harmonic generation [15] and on-chip squeezing [12], where device statistics and variations make unbiased coupling calibration essential. Beyond the two processes highlighted in this work, the same strategy can be applied to a broad class of nonlinear interactions such as SPDC [3, 4, 5], FWM [6, 7, 8, 9, 10, 11, 12, 13], OPO [19, 20, 21, 22] and OPA [23, 3, 16], providing a unified approach for coupling-resolved benchmarking as integrated nonlinear photonics continues to scale.
Conceptually, the complementary SHG and SPDC mechanisms used here reflect a broader principle: nonlinear processes introduce measurement asymmetries that break the intrinsic $\eta_{1}\eta_{2}$ degeneracy that limits purely linear transmission-based calibration methods. In practice, any bidirectional or multi-process approach naturally exhibits small forward and backward differences, for example, in phase matching or mode overlap. These effects can similarly be utilized for BNOT, given that the essential requirement is a controlled asymmetry in the forward model that distinguishes the two interfaces in a physically meaning way. When this asymmetry is taken into account, the joint efficiency inference converges reliably to the true coupling values.
Looking forward, a promising direction is to integrate BNOT-style tomography into adaptive control and self-calibrating photonic architectures. Real-time forward and backward measurements, combined with Bayesian [47, 48, 49, 50] or machine-learning estimators [51], may enable automatic compensation of coupling losses across large photonic arrays and continuous tracking of drift or degradation. Such closed-loop implementations have the potential to transform coupling calibration from a post-measurement diagnostic into an active subsystem of photonic hardware, supporting reproducible, system-level benchmarks as nonlinear PICs transition from laboratory demonstrations to scalable quantum and classical technologies.
Acknowledgements.
The authors thank Mengjie Yu, Avik Dutt, and Sri Krishna Vadlamani for fruitful discussions. This work was supported by funding from the DARPA INSPIRED program.
Appendix A Analytical formula derivation
In this section, we analytically derive the squeezing and SHG dynamics generated in a PPLN device. Let $\hat{a}_{\omega}$ and $\hat{a}_{2\omega}$ denote the field operators at frequencies $\omega$ and $2\omega$, respectively. Both nonlinear interactions arise from the same three–wave mixing Hamiltonian at PPLN position $z$, given by
$\hat{\mathcal{H}}=\frac{\hbar g}{2i}\left(\hat{a}_{\omega}^{2}\,\hat{a}_{2\omega}^{\dagger}\,e^{i\Delta kz}-\hat{a}_{\omega}^{\dagger\,2}\,\hat{a}_{2\omega}\,e^{-i\Delta kz}\right),$ (5)
where $g$ is the nonlinear coupling strength, $\Delta k=2k_{\omega}-k_{2\omega}$ is the wave-vector mismatch.
Thus, the temporal evolution of the operators are governed by the Heisenberg–Langevin equations
$\frac{d\,\hat{a}_{2\omega}}{dz}$ $=-\frac{g}{2}\,\hat{a}^{2}_{\omega}\,e^{i\Delta kz}-\frac{\Gamma_{2\omega}}{2}\,\hat{a}_{2\omega}+\sqrt{\Gamma_{2\omega}}\,\hat{v}_{2\omega}$ (6)
$\frac{d\,\hat{a}_{\omega}}{dz}$ $=g\,\hat{a}_{\omega}^{\dagger}\,\hat{a}_{2\omega}\,e^{-i\Delta kz}-\frac{\Gamma_{\omega}}{2}\,\hat{a}_{\omega}+\sqrt{\Gamma_{\omega}}\,\hat{v}_{\omega},$
where $\Gamma_{2\omega}$ and $\Gamma_{\omega}$ are the PPLN waveguide propagation losses at the corresponding modes, $\hat{a}_{2\omega}$ and $\hat{a}_{\omega}$ are the field operators, $\hat{v}_{2\omega}$ and $\hat{v}_{\omega}$ are the vacuum noise operators. The field and vacuum operators satisfy the commutation relations:
$\left[\hat{a}_{\nu},\hat{a}^{\dagger}_{\mu}\right]=\left[\hat{v}_{\nu},\hat{v}^{\dagger}_{\mu}\right]=\delta_{\nu\mu}\quad,\quad\langle\hat{a}^{\dagger}_{\nu}\hat{a}_{\mu}\rangle=\langle\hat{v}^{\dagger}_{\nu}\hat{v}_{\mu}\rangle=0,$ (7)
for $\nu,\,\mu\in\{2\omega,\omega\}$. Although both squeezing and SHG processes originate from the same interaction Hamiltonian, the relevant approximations and boundary conditions differ. In the following subsections, we solve the Heisenberg–Langevin equation of SHG and squeezing.
A.1 Second-harmonic generation
In the SHG process, the pump field at frequency $\omega$ is treated as a strong, undepleted classical drive such that $\langle\hat{a}_{\omega}\rangle\approx A_{\omega}$ remains constant along the waveguide and the non-linear coupling strength of SHG is
$g_{\text{SHG}}=\frac{2\omega\,d_{\text{eff}}}{n_{2\omega}\,c}\sqrt{\frac{\hbar\omega\,S_{2\omega}}{\epsilon_{0}\,S^{2}_{\omega}\,L}},$
where $d_{\text{eff}}=2d_{33}/\pi$ is the effective nonlinear coefficient under quasi-phase-matching, $d_{33}$ is the nonlinear optical coefficient, $c$ is the speed of light, $\epsilon_{0}$ denotes the vacuum permittivity, $n_{2\omega}$ is the refractive index of LiN at $2\omega$, $S_{2\omega}$ and $S_{\omega}$ are the effective mode areas at the corresponding modes, and $L$ is the poling length of PPLN.
Under this approximation, the second-harmonic mode $\hat{a}_{2\omega}$ obeys a driven linear equation, and its mean-field solution at position $z=L$ is
$\langle\hat{a}_{2\omega}(L)\rangle$ $\approx-\frac{g_{\text{SHG}}\,A_{\omega}^{2}}{2}\,\int_{0}^{L}\,dz^{\prime}\,e^{\left(i\Delta k+\frac{\Gamma_{2\omega}}{2}\right)\,(z^{\prime}-L)}$ (8)
$=-\frac{g_{\text{SHG}}\,A_{\omega}^{2}}{2}\,\left(\frac{1-e^{-\left(i\Delta k+\frac{\Gamma_{2\omega}}{2}\right)\,L}}{i\Delta k+\Gamma_{2\omega}/2}\right).$
where $\Delta k$ is the wave-vector mismatch, and $\Gamma_{2\omega}$ is the loss of the generated second-harmonic field.
$|\langle\hat{a}_{2\omega}(L)\rangle|^{2}=g_{\text{SHG}}^{2}\,A_{\omega}^{4}\,\left(\frac{1+e^{-\Gamma_{2\omega}L}-2e^{-\frac{\Gamma_{2\omega}L}{2}}\cos{\left(\Delta kL\right)}}{4\Delta k^{2}+\Gamma^{2}_{2\omega}}\right).$ (9)
To quantify the overall frequency-conversion performance, we express the generated second-harmonic signal in terms of optical power. For a mode at frequency $\Omega\in\{\omega,2\omega\}$, the optical power is related to its electric-field amplitude $E_{\Omega}$ by
$P_{\Omega}=\frac{n_{\Omega}\,c\,\epsilon_{0}}{2}\,|E_{\Omega}|^{2}S_{\Omega}.$ (10)
The fields $E_{2\omega}$ and $E_{\omega}$ are obtained by normalizing the dimensionless amplitudes $\langle\hat{a}_{2\omega}(L)\rangle$ and $A_{\omega}$ to the corresponding single-photon electric fields. In particular, they are related through
$E_{2\omega}=\langle\hat{a}_{2\omega}(L)\rangle\sqrt{\frac{4\hbar\omega}{n_{2\omega}^{2}\epsilon_{0}\,S_{2\omega}L}}\quad,\;E_{\omega}=A_{\omega}\sqrt{\frac{2\hbar\omega}{n_{\omega}^{2}\epsilon_{0}\,S_{\omega}L}}.$ (11)
In this way, $\left|\langle\hat{a}_{2\omega}(L)\rangle\right|^{2}$ and $\left|A_{\omega}\right|^{2}$ directly represent the mean photon numbers at $2\omega$ and $\omega$, while $E_{2\omega}$ and $E_{\omega}$ carry the proper physical units required in the power expression of Eq. (10).
Using this relation, the on-chip second-harmonic power $P^{(\text{ON})}_{2\omega}$ referenced to the off-chip pump power $P_{\omega}^{(\text{OFF})}$ leads to the on-chip SHG efficiency [17]:
$\mathcal{E}^{(\varepsilon_{2},\,\Delta k)}_{\text{ON}}(x_{2})$ $\equiv\frac{P^{(\text{ON})}_{2\omega}}{\left(P_{\omega}^{(\text{OFF})}\,L\right)^{2}}=\frac{2\,n_{2\omega}\,S_{2\omega}}{n^{2}_{\omega}\,\epsilon_{0}\,c\,L^{2}\,S^{2}_{\omega}}\left|\frac{E_{2\omega}}{E_{\omega}^{(\text{OFF})\,2}}\right|^{2}$ (12)
$=\beta\,x_{2}^{2}\,\varepsilon_{2}^{2}\,\left(\frac{1+e^{-\Gamma_{2\omega}L}-2e^{-\frac{\Gamma_{2\omega}L}{2}}\cos{\left(\Delta kL\right)}}{\Delta k^{2}+\Gamma^{2}_{2\omega}/4}\right),$
where
$\beta=\frac{8\pi^{2}d^{2}_{\text{eff}}}{n_{2\omega}\,n_{\omega}^{2}\,c\,\epsilon_{0}\,L^{2}\,\lambda_{\omega}^{2}}\frac{S_{2\omega}}{S_{\omega}^{2}},$
i.e., $\lambda_{\omega}=2\pi c/\omega$. Here, $E_{\omega}^{(\text{OFF})}$ denotes the off-chip pump electric field, and $n_{\omega}$ is the refractive index of LiN at $\omega$. After coupling into the PPLN waveguide, it becomes $E_{\omega}=x_{2}\,\varepsilon_{2}\,E_{\omega}^{(\text{OFF})}$ with tunable coupling efficiency $x_{2}$ and pump-power fluctuation rate $\varepsilon_{2}$.
A.2 Squeezed light
Squeezed-light generation in the PPLN waveguide arises from the SPDC process in which the $2\omega$ pump is converted into photon pairs at $\omega$. Under strong pumping, the $2\omega$ mode is treated as a classical undepleted drive, $\langle\hat{a}_{2\omega}\rangle\approx A_{2\omega}$, which leads to
$\frac{d\,\hat{a}_{\omega}}{dz}=r\,e^{-i\Delta kz}-\frac{\Gamma_{\omega}}{2}\,\hat{a}_{\omega}+\sqrt{\Gamma_{\omega}}\,\hat{v}_{\omega},$ (13)
where $r\equiv g_{\text{SQZ}}\,A_{2\omega}\,L$ denotes the on-chip squeezing level, coupling strength
$g_{\text{SQZ}}=\frac{\omega\,d_{\text{eff}}}{n_{\omega}\,c}\sqrt{\frac{4\hbar\omega}{n_{2\omega}^{2}\,\epsilon_{0}\,S_{2\omega}\,L}}.$
Unlike SHG, Eq. (13) admits an additional simplification. In a single-pass waveguide, both the nonlinear gain and the propagation loss accumulate gradually, and over an infinitesimal segment $dz$ they act on the same mode in a way that does not meaningfully interfere. To the first order $\mathcal{O}(dz)$, it makes no physical difference whether a small amount of loss occurs before or after a small amount of squeezing. Consequently, the entire distributed evolution can be modeled as an ideal Bogoliubov transformation followed by a single effective loss channel.
Under this approximation, the squeezing unitary operation is described by
$\hat{U}(L)$ $=\exp{\left[\frac{r}{2}\,\left(\hat{a}^{2}_{\omega}\,\left(\int_{0}^{L}\,dz\,e^{i\,\Delta k\,z}\right)-\hat{a}^{\dagger\,2}_{\omega}\,\left(\int_{0}^{L}\,dz\,e^{-i\,\Delta k\,z}\right)\right)\right]}$ (14)
$=\exp{\left[\frac{r}{2}\,\left(\hat{a}^{2}_{\omega}\,\left(\frac{e^{i\,\Delta k\,L}-1}{i\,\Delta\,k}\right)-\hat{a}^{\dagger\,2}_{\omega}\,\left(\frac{e^{-i\,\Delta k\,L}-1}{-i\,\Delta\,k}\right)\right)\right]}$
$=\exp{\left[\frac{r^{\prime}}{2}\,\left(\hat{b}^{2}_{\omega}-\hat{b}^{\dagger\,2}_{\omega}\right)\right]},$
where $\phi=\text{Arg}\left(\frac{e^{i\,\Delta k\,L}-1}{i\,\Delta\,k}\right)$, $r^{\prime}=r\left|\text{sinc}\left(\frac{\Delta k\,L}{2}\right)\right|$ and $\hat{b}_{\omega}=\hat{a}_{\omega}\,e^{i\phi/2}$. Loss is treated as a single amplitude-damping channel with total transmission $\eta(L)=e^{-\Gamma_{\omega}L}$. The output field operator in the rotating frame ($\hat{b}_{\omega}$) is therefore
$\hat{b}_{\omega}(L)$ $\approx\sqrt{\eta(L)}\,\left(\hat{U}^{\dagger}(L)\,\hat{b}_{\omega}\,\hat{U}(L)\right)+\sqrt{1-\eta(L)}\,\hat{v}_{\omega}$ (15)
$=e^{-\frac{\Gamma_{\omega}\,L}{2}}\,\left(\hat{b}_{\omega}\,\cosh{r^{\prime}}-\hat{b}^{\dagger}_{\omega}\,\sinh{r^{\prime}}\right)+\sqrt{1-e^{-\Gamma_{\omega}\,L}}\,\hat{v}_{\omega}.$
Evaluating the field at the output $z=L$ gives
$\hat{q}_{\omega}(L)$ $=e^{-\frac{\Gamma_{\omega}\,L}{2}}\,\left(\hat{b}_{\omega}+\hat{b}^{\dagger}_{\omega}\right)\,e^{-r^{\prime}}+\sqrt{1-e^{-\Gamma_{\omega}\,L}}\,\left(\hat{v}_{\omega}+\hat{v}^{\dagger}_{\omega}\right)$ (16)
$\hat{p}_{\omega}(L)$ $=e^{-\frac{\Gamma_{\omega}\,L}{2}}\,\left(\frac{\hat{b}_{\omega}-\hat{b}^{\dagger}_{\omega}}{i}\right)\,e^{r^{\prime}}+\sqrt{1-e^{-\Gamma_{\omega}\,L}}\,\left(\frac{\hat{v}_{\omega}-\hat{v}^{\dagger}_{\omega}}{i}\right),$
using the conventions $\hat{q}_{\omega}=\hat{a}_{\omega}+\hat{a}_{\omega}^{\dagger}$ and $\hat{p}_{\omega}=(\hat{a}_{\omega}-\hat{a}_{\omega}^{\dagger})/i$. Since the input and vacuum modes have zero mean, $\langle\hat{q}_{\omega}(L)\rangle=\langle\hat{p}_{\omega}(L)\rangle=0$.
The quadrature variances follow directly [46]:
$\Delta q^{2}_{\omega}(L)$ $=e^{-\Gamma_{\omega}\,L}\,e^{-2\,r^{\prime}}+1-e^{-\Gamma_{\omega}\,L}<1\quad\quad\text{(squeezing)}$ (17)
$\Delta p^{2}_{\omega}(L)$ $=e^{-\Gamma_{\omega}\,L}\,e^{2\,r^{\prime}}+1-e^{-\Gamma_{\omega}\,L}>1\quad\text{(anti-squeezing)}.$
To relate the squeezing level of the field to the pump power, we express the pump amplitude as
$A_{2\omega}$ $=\sqrt{\frac{2\,x_{1}\,\varepsilon_{1}\,P^{(0)}_{2\omega}}{n_{2\omega}\,c\,\epsilon_{0}\,S_{2\omega}}\left(\frac{n_{2\omega}^{2}\,\epsilon_{0}\,S_{2\omega}\,L}{4\,\hbar\,\omega}\right)}=\sqrt{\frac{n_{2\omega}\,L\,x_{1}\,\varepsilon_{1}\,P^{(0)}_{2\omega}}{2\,\hbar\,\omega\,c}},$ (18)
where $P^{(0)}_{2\omega}$ is the amplitude of the off-chip pump power, $x_{1}$ is the tunable off-chip coupling efficiency, and $\varepsilon_{1}$ is the pump fluctuation rate. As a result, the resulting on-chip squeezing level is
$\mathcal{S}^{(\varepsilon_{1},\,\Delta k)}_{\text{ON}}$ $(x_{1})=$ (19)
$-\log\left(e^{-\Gamma_{\omega}L}e^{-2\sqrt{x_{1}\varepsilon_{1}\alpha P^{(0)}_{2\omega}}\,\left|\text{sinc}\left(\frac{\Delta kL}{2}\right)\right|}+1-e^{-\Gamma_{\omega}L}\right),$
where $\alpha=\sqrt{n_{2\omega}\,L/2\hbar\omega c}$.
Appendix B Efficiency thresholds for off-chip squeezing
Figure 5: (a) On-chip squeezing versus $\eta_{1}^{(2\omega)}$. (b) Off-chip squeezing versus $\eta_{1}^{(2\omega)}$ and $\eta_{2}^{(\omega)}$. The orange horizontal line in (a) and curve in (b) denote the on-chip and off-chip $15$-dB squeezing threshold.
Fig. 5 summarizes the dependence of on- and off-chip squeezing on interface efficiencies. Fig. 5(a) shows the on-chip squeezing $\mathcal{S}_{\text{ON}}$ as a function of the coupling efficiency $\eta^{(2\omega)}_{1}$, while Fig. 5(b) maps the off-chip squeezing $\mathcal{S}_{\text{OFF}}$ as a joint function of $\eta^{(2\omega)}_{1}$ and $\eta^{(\omega)}_{2}$. The orange lines in both panels mark the squeezing contour $15\penalty 10000\ \text{dB}$, a benchmark frequently cited in advanced quantum technologies.
On-chip squeezing levels approaching $15\penalty 10000\ \text{dB}$ are particularly relevant for applications such as CV quantum error correction and precision quantum metrology [26, 39]. As shown in Fig. 5(a), achieving this on-chip benchmark requires an input coupling efficiency $\eta^{(2\omega)}_{1}\gtrsim 0.65$, which is attainable using current photonic integration technologies. For example, edge-coupled LiN modulators have demonstrated per-interface coupling efficiencies below $0.5\penalty 10000\ \text{dB}$ (corresponding to efficiencies $90\penalty 10000\ \%$) at $1550\penalty 10000\ \text{nm}$ [52], indicating that such coupling efficiencies are experimentally feasible.
Although high on-chip squeezing is feasible, maintaining this performance off chip imposes stricter requirements. For example, to retain the off-chip squeezing level $\mathcal{S}_{\text{OFF}}\gtrsim 15\penalty 10000\ \text{dB}$, the output interface efficiency must satisfy $\eta^{(\omega)}_{2}\gtrsim 0.97$, illustrated in Fig. 5(b).
Simulation parameters $d_{33}$ $19.5$ pm/V Thin-film LiN [17] nonlinear coefficient $L$ 11 mm  Poling length of PPLN $\omega$ $193$ THz Pumping frequency (or $2\omega$)  (or $386$ THz)   at $1550$ nm (or $775$ nm) $n_{\omega}$ $2.14$  Refractive index of LiN (or $n_{2\omega}$) (or $2.2$)  at $\omega$ (or $2\omega$) $\mathcal{S}_{\omega}$ 11 µm2 Mode area of the field (or $\mathcal{S}_{2\omega}$)  (or $0.8$ µm2) in PPLN waveguide $\Gamma_{\omega}$ 55 dB/m [53] Propagation loss (or $\Gamma_{2\omega}$) (or $10$ dB/m) $\sigma_{\text{sqz}}$ $0.1\penalty 10000\ \text{dB}$  Squeezing (or SHG) (or $\sigma_{\text{shg}}$)    (or $20\penalty 10000\ \text{\%/W/cm${}^{2}$}$)   measurement noise STD Off-chip average $P_{2\omega}$ $100$ mW pump power in the squeezing process Half range of power $\delta_{1}$ $0.5\penalty 10000\ \%$ fluctuation rate in (or $\delta_{2}$) squeezing (or SHG) $\sigma_{\Delta k}$ 0.36 m-1 Wave-vector mismatch STD Linear transmission $\sigma_{\text{C}}$ $7\penalty 10000\ \%$ measurement STD with conventional approach
Table 1: Monte Carlo simulation parameters of the PPLN waveguide in squeezing and SHG processes. STD: standard deviation.